OpenStudy (anonymous):

Given real numbers m,n,p, and q with n,p and q being rational, and such that m/3 + 3/2n = 4p - q/2, show that m must be a rational number too.

OpenStudy (anonymous):

\[\frac{m}{3} + \frac{3n}{2} = 4 p - \frac{q}{2}\]

OpenStudy (anonymous):

thats the equation, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so you just need to prove that you can express m as a fraction

OpenStudy (anonymous):

write n, p, and q in fraction form and then isolate m. As long as m is equal to a fraction, then it must be also a rational number

OpenStudy (anonymous):

Ah, I think I got it. Thanks.